The moons mass is 7.4×1022 kg, and its radius is 1.7×106 m. Use newton theory of gravity to find the weight of a 1kg object lying on the moons surface.


1
Expert's answer
2021-02-17T10:39:59-0500

The gravitational force of attraction on a 1.0 kg object (or its weight) at the surface of the Moon can be calculated as follows:


"F_M=\dfrac{GM_Mm}{R_M^2},""F_M=\dfrac{6.67\cdot10^{-11}\ \dfrac{Nm^2}{kg^2}\cdot7.35\cdot10^{22}\ kg\cdot1\ kg}{(1.738\cdot10^6\ m)^2}=1.62\ N."